OpenStudy (anonymous):

Is the point (-3,4) a solution of the given inequality? 3x - 7y > 14 (Step by step please) Thanks

OpenStudy (anonymous):

the question is asking this : is \[3\times (-3)-7\times 4>14\] or not ?

jaynator495 (jaynator495):

3x-7y>14 3 x -3 - 7 x 4 > 14 3 x -3 = 9 7 x 4 = 28 9 - 28 = 19 19 > 14 ? so that is a solution to the given inequality because 19 is greather then 14, theres your step by step awnser :) hope this helped, -Jaynator495

jaynator495 (jaynator495):

@holly01